# Rasterized Steered Mixture of Experts for Efficient 2D Image Regression

### v1

@@ I Introduction
Image regression estimates a continuous function that maps image coordinates to pixel values. It enables smooth, high-fidelity representations of visual content and often employs neural networks or kernel-based models to enhance generalization and preserve fine details [1, 2, 3, 4, 5].
@@
In practice, image regression frequently involves reconstructing missing or corrupted data, casting it as an inverse problem. This problem is inherently ill-posed, requiring regularization or learned priors for stable recovery [6, 7]. While classical approaches rely on handcrafted priors, modern methods employ deep networks to learn implicit structures [8], or leverage powerful techniques like plug-and-play priors [9] and score-based diffusion models [10].
@@
Sparsity offers a powerful inductive bias for addressing ill-posedness. Sparse models reduce complexity, enhance interpretability, and support better generalization, especially under limited or noisy observations. Classical frameworks such as compressed sensing [11] and dictionary learning [12] exploit sparsity for signal recovery, while recent advances embed similar ideas within deep architectures via sparse coding networks [13] and unrolled optimization schemes [14, 15].
@@
Recent works have advanced kernel-based methods for image regression and restoration by extending classical Gaussian models to deep or hybrid formulations. Quan et al. [16] proposed a deep Gaussian kernel mixture for single-image defocus deblurring, demonstrating the ability of Gaussian mixtures to handle complex inverse problems while preserving fine structures. Li et al. [17] introduced an adaptive segmentation-based initialization strategy for steered mixture of experts and kernel regression, highlighting the importance of structured initialization in accelerating convergence and improving sparsity. Complementing these practical advances, Medvedev et al. [18] analyzed the overfitting behavior and generalization bounds of ridgeless Gaussian kernel regression, offering theoretical insights relevant to sparse kernel-based regression frameworks. These studies underscore the continued relevance of Gaussian kernels and their variants for efficient and high-quality image regression, aligning closely with the principles motivating our SMoE approach.
@@
Motivated by the demands of sparse image regression, we focus on the Steered Mixture of Experts (SMoE) framework, a Gaussian-based model designed to efficiently handle sparse image regression through expert blending. SMoE’s effectiveness has been demonstrated in 2D sparse image representation [19], 2D and 3D image and video compression [20], and 4D/5D light field representation and coding [21]. While steered Gaussian kernels dominate these models, strong performance has also been reported using steered Epanechnikov kernels, particularly for image and light field data [22]. SMoE has also shown promise in classic low-level vision applications: its edge-aware formulation has recently been leveraged for image denoising and restoration, yielding competitive results in a domain that has seen sustained attention over the past two decades [23, 24]. Overall, SMoE presents a versatile and principled approach used for denoising, super-resolution, and high-dimensional data representation.
@@
SMoE combines sparsity with edge-awareness, making it especially well-suited for inverse problems. At the core of the SMoE model is an edge-aware kernel representation, which adapts to the local structures in images, preserving crucial details while performing operations such as sparse representation for regression and compression, noise reduction, and resolution enhancement. Sparsity and edge-preservation in SMoE regression models are achieved using a gating network, i.e. with normalized steered Gaussian kernels. Well-known kernel regression frameworks, such as Radial Basis Function networks in Gaussian Splatting (GS-RBF) [4] or Takeda Kernel Regression [25], model pixel values as “weighted sum of kernels”. In contrast, the SMoE gating network models pixels as “weighted sum of gating functions” using strategies similar to normalized RBF [26] networks. This ensures the sparsity of the SMoE model – few kernels are sufficient to represent complex textures. An optimization process ensures that SMoE gating functions align with edges in images easily and efficiently. Sharp edges and smooth transitions in images are represented with sparse SMoE models using few kernels. By comparison, RBF regression frameworks, on the other hand, usually require many kernel functions for edge reconstruction.
@@
Despite SMoE’s sparsity and edge-awareness, optimizing these models remains a central challenge. The high parameter count and non-convex objective landscape make gradient descent slow and computationally expensive. For each image, a unique set of optimal parameters must be identified. As with RBF network optimization, SMoE optimization typically requires minimizing highly non-convex objective functions with thousands of parameters. This can result in excessive runtimes, rendering many SMoE approaches impractical for real-world deployment. There are primarily two strategies for implementing SMoE models for regression: (a) block-wise SMoE [27, 28], which optimizes kernels locally using either autoencoders or iterative solvers, and (b) global SMoE [19, 29, 30, 20], which jointly optimizes all kernels over the full image. While block-wise SMoE offers speed, its local view can compromise quality. Global SMoE delivers greater sparsity and fidelity, but at the cost of significant computational effort.
@@
Gaussian Splatting (GS) [4, 5], a recent method originally designed for 3D scene representation and rendering using volumetric Gaussian kernels—and conceptually related to earlier surface splatting techniques [31], shares several similarities with SMoE. Both methods splat steered Gaussians into the pixel domain and use gradient descent to optimize Gaussian kernels for effective data representation. GS leverages localized block-rasterization to accelerate training, primarily targeting 3D applications. The key distinction between GS and SMoE lies in the GS-RBF regression model versus the SMoE normalized gating network for regression. SMoE regression can be seen as an extended, more powerful strategy compared to GS-RBF with normalization using the same kernel representation.
@@
First attempts to adopt the rasterized GS strategy to 2D image regression appeared in [5]. Building decisively on this foundation, our work is the first to integrate GS-inspired rasterization directly into the SMoE framework, combining GS’s computational speed with SMoE’s superior sparsity and edge-aware reconstruction. The purpose of this paper is to take advantage of both GS (fast) and SMoE (sparse and high-quality reconstruction) approaches to arrive at a fast and sparse high-quality SMoE regression method. To this end, we seek to adopt the rasterization approach of GS to optimize SMoE parameters, overcoming the critical bottleneck of slow, iterative gating network optimization and enabling drastic run-time improvements.
@@
  • We introduce R-SMoE, a rasterized SMoE training framework that leverages tile-based rasterization inspired by Gaussian Splatting. Compared to non-rasterized SMoE regression, R-SMoE accelerates model training and rendering times by orders of magnitude while maintaining high fidelity. Compared to “GaussianImage” [5], which employs GS-RBF rasterization, our approach requires significantly fewer computational resources for both training and rendering.
@@
Figure 1: Illustration of the edge-aware Steered Mixture of Experts (SMoE) model applied to image compression and denoising at 0.43 bits per pixel (bpp). Figure adapted from [28].
@@ II Steered Mixture of Experts (SMoE)
The SMoE image model describes an edge-aware, parametric, continuous nonlinear regression function. Fig. 1 (from [28]) illustrates the soft-gating concept of the kernel model for image compression and denoising tasks.
@@ II-A Theory
The Steered Mixture of Experts (SMoE) model extends traditional kernel regression methods, such as Radial Basis Function (RBF) [32] networks used in Gaussian Splatting, by introducing a more sophisticated mechanism for image reconstruction. Although Gaussian Splatting also uses a “weighted sum of Gaussians,” which is conceptually similar to RBF regression, the interpretation differs: in RBF, the weights represent the importance of each Gaussian basis, whereas in Gaussian Splatting, the weights correspond to the color (appearance) of the Gaussian. Thus, while the regression form is similar, the modeling objectives are fundamentally different. RBF networks define the regression function $y_{p}(x)$ based on a weighted sum of $L$ (steered 2D Gaussian) kernels $K_{j}(x)$:
@@
The $\pi_{i}$ values provides additional weights to each kernel. Expert functions $m_{j}(x)$ can take various functional forms, including constant, linear, quadratic, and basis functions such as DCT and wavelet bases [33]. For our work, we utilize simple constant experts $m_{j}(x)=m_{j}$, which have demonstrated excellent performance in previous studies [19, 27]. The irregularly shaped 2D softmax gating functions $w_{j}(x)$ are derived based on the position and parameters of the $L$ kernels, allowing for precise modeling of both smooth transitions and sharp edges in images.
@@ III-A Deep-Learning-Based Optimization
Recently, several methods have been proposed to replace traditional gradient descent (GD) optimization with deep-learning approaches. These methods train neural networks to directly predict kernel parameters. For instance, Fleig et al. [28, 24] trained an auto-encoder network to predict kernel parameters with a SMoE regressor according to Eq. (3) as the decoder. This results in drastically faster optimization, replacing the cumbersome GD process. However, these approaches are currently applicable for a maximum of 16 kernels on small blocks of 8x8 or 16x16 pixels. Currently, most approaches use a fixed number of kernels for each block, which provides limited adaptation to content. Additionally, because kernel parameter prediction is based on small, isolated blocks, independent predictions across different blocks can introduce blocking effects, where discontinuities or artifacts may appear between blocks.
@@ III-B Global Optimization
Global optimization [19, 20] represents the forefront of optimization strategies for kernel regression and SMoE models. This method optimizes the model by considering the entire set of kernels simultaneously, rather than treating them individually or in isolated groups. The core idea is to ensure that all kernels in the model contribute to the reconstruction of each pixel, thereby capturing the holistic structure of the image. This strategy is also common in RBF type regression [34]. Global optimization thus involves using all kernels to jointly reconstruct each pixel. Strategies exist to constrain the impact of kernels to the immediate neighborhood [35] to reduce the complexity of reconstruction.
@@ III-C Rasterized optimization
Recently, rasterization [36, 37] has become an increasingly valuable tool for enhancing gradient descent (GD) optimization in image processing tasks. One notable example is the GS for 2D images (GaussianImage [5]), which leverages rasterization to boost the efficiency of GD optimization in Radial Basis Function (RBF) [32] networks. Adopted from 3D GS [4], this approach is similar to deep-learning-based optimization, being block-based in nature. Each image is divided into adjacent, non-overlapping blocks of 16x16 pixels. A very large number of 2D kernels are initialized over the image domain. A small subset of the kernels is identified as relevant for regressing a block prior to GD optimization, which allows fast and accurate GD optimization and block reconstruction. In this paper, we adapt the GS rasterized optimization strategy to SMoE regression, resulting in fast, sparse, and high-quality reconstruction.
@@ III-D Hybrid EM-Based Parallel Rendering
In addition to gradient-descent-based rasterized optimization, prior work by Avramelos et al. [38] introduced a block-parallel rendering method for SMoE models, designed for real-time applications. Their method relies on a two-phase pipeline: a global Expectation-Maximization (EM) optimization followed by block-level parallel rendering. During optimization, Gaussian kernel parameters—position, orientation, and scale—are learned globally using EM. The rendering stage then employs a fixed-radius search around each pixel to determine which kernels to evaluate, enabling efficient per-pixel computation.
@@ IV Rasterized SMoE
The pipeline for the proposed rasterized SMoE is illustrated in Fig. 3 (a). This pipeline mirrors that of [5], except that the RBF regression is replaced with the SMoE gating framework (c.f. Equ. (2) and (4)).
@@ IV-B Geometry shader
For each 16x16 pixel block, each Gaussian kernel of the set $\mathcal{K}$ is given a confidence ellipse corresponding to a 99% confidence interval of its 2D Gaussian distribution. This ellipse approximates the region in which the kernel has significant influence, with its major and minor axes derived from the kernel’s covariance matrix. For the mathematical intricacies—including axis derivation via eigen-decomposition—see [39].
@@ V Rasterized SMoE for Denoising
While previous SMoE methods have employed block-based representations for denoising, either through gradient descent (GD) [23] or deep learning (DL) [24] optimization, these approaches are limited by their block constraints. R-SMoE promises a more efficient and reliable representation, combining the advantages of global representation with the flexibility of local processing. R-SMoE does not confine the influence of kernels to a small block, allowing each kernel to contribute beyond its immediate neighborhood.
One significant challenge when applying GS for 2D images and SMoE models to denoising tasks is that, when applied to smaller blocks of 16x16 pixels, these models attempt to accurately reconstruct not only the signal but also part of the noise. A promising strategy is then to employ overlapping GS blocks or R-SMoE blocks and reconstruct the pixels using multiple models in a multi-model approach [23]. This yields a natural extension: multi-model inference in R-SMoE (MM-RSMoE), which we explore for denoising.

### v2

@@ I Introduction
Image regression estimates a continuous function that maps image coordinates to pixel values. It enables smooth, high-fidelity representations of visual content and often employs neural networks or kernel-based models to enhance generalization and preserve fine details [mildenhall2021nerf, salehi2022deep, Li_2023_CVPR, kerbl_3d_2023, zhang2024gaussianimage].
@@
In practice, image regression frequently involves reconstructing missing or corrupted data, casting it as an inverse problem. This problem is inherently ill-posed, requiring regularization or learned priors for stable recovery [bertero2021introduction, hasanouglu2021introduction]. While classical approaches rely on handcrafted priors, modern methods employ deep networks to learn implicit structures [genzel2022solving], or leverage powerful techniques like plug-and-play priors [zhang2021plug] and score-based diffusion models [croitoru2023diffusion].
@@
Sparsity offers a powerful inductive bias for addressing ill-posedness. Sparse models reduce complexity, enhance interpretability, and support better generalization, especially under limited or noisy observations. Classical frameworks such as compressed sensing [Baraniuk2010compressivesensing] and dictionary learning [Mohamad2023dictionary] exploit sparsity for signal recovery, while recent advances embed similar ideas within deep architectures via sparse coding networks [Yuli2022structured] and unrolled optimization schemes [Kofler2023unrolling, Zhao2022imagefusionunrolling].
@@
Recent works have advanced kernel-based methods for image regression and restoration by extending classical Gaussian models to deep or hybrid formulations. Quan et al. [quan2024gaussiandefocus] proposed a deep Gaussian kernel mixture for single-image defocus deblurring, demonstrating the ability of Gaussian mixtures to handle complex inverse problems while preserving fine structures. Li et al. [li2025adptsmoe] introduced an adaptive segmentation-based initialization strategy for steered mixture of experts and kernel regression, highlighting the importance of structured initialization in accelerating convergence and improving sparsity. Complementing these practical advances, Medvedev et al. [medvedev2025gaussianridgeless] analyzed the overfitting behavior and generalization bounds of ridgeless Gaussian kernel regression, offering theoretical insights relevant to sparse kernel-based regression frameworks. These studies underscore the continued relevance of Gaussian kernels and their variants for efficient and high-quality image regression, aligning closely with the principles motivating our SMoE approach.
@@
Motivated by the demands of sparse image regression, we focus on the Steered Mixture of Experts (SMoE) framework, a Gaussian-based model designed to efficiently handle sparse image regression through expert blending. SMoE’s effectiveness has been demonstrated in 2D sparse image representation [bochinski_regularized_2018], 2D and 3D image and video compression [jongebloed_sparse_2022], and 4D/5D light field representation and coding [verhack_steered_2020]. While steered Gaussian kernels dominate these models, strong performance has also been reported using steered Epanechnikov kernels, particularly for image and light field data [liu_4d_2022]. SMoE has also shown promise in classic low-level vision applications: its edge-aware formulation has recently been leveraged for image denoising and restoration, yielding competitive results in a domain that has seen sustained attention over the past two decades [ozkan_steered_2023, fleig_steered_2023]. Overall, SMoE presents a versatile and principled approach used for denoising, super-resolution, and high-dimensional data representation.
@@
SMoE combines sparsity with edge-awareness, making it especially well-suited for inverse problems. At the core of the SMoE model is an edge-aware kernel representation, which adapts to the local structures in images, preserving crucial details while performing operations such as sparse representation for regression and compression, noise reduction, and resolution enhancement. Sparsity and edge-preservation in SMoE regression models are achieved using a gating network, i.e. with normalized steered Gaussian kernels. Well-known kernel regression frameworks, such as Radial Basis Function networks in Gaussian Splatting (GS-RBF) [kerbl_3d_2023] or Takeda Kernel Regression [takeda2007kernelregression], model pixel values as “weighted sum of kernels”. In contrast, the SMoE gating network models pixels as “weighted sum of gating functions” using strategies similar to normalized RBF [Heimes1998nrbf] networks. This ensures the sparsity of the SMoE model – few kernels are sufficient to represent complex textures. An optimization process ensures that SMoE gating functions align with edges in images easily and efficiently. Sharp edges and smooth transitions in images are represented with sparse SMoE models using few kernels. By comparison, RBF regression frameworks, on the other hand, usually require many kernel functions for edge reconstruction.
@@
Despite SMoE’s sparsity and edge-awareness, optimizing these models remains a central challenge. The high parameter count and non-convex objective landscape make gradient descent slow and computationally expensive. For each image, a unique set of optimal parameters must be identified. As with RBF network optimization, SMoE optimization typically requires minimizing highly non-convex objective functions with thousands of parameters. This can result in excessive runtimes, rendering many SMoE approaches impractical for real-world deployment. There are primarily two strategies for implementing SMoE models for regression: (a) block-wise SMoE [tok_mse_2018, fleig_edge-aware_2022], which optimizes kernels locally using either autoencoders or iterative solvers, and (b) global SMoE [bochinski_regularized_2018, jongebloed_hierarchical_2018, jongebloed_quantized_2019, jongebloed_sparse_2022], which jointly optimizes all kernels over the full image. While block-wise SMoE offers speed, its local view can compromise quality. Global SMoE delivers greater sparsity and fidelity, but at the cost of significant computational effort.
@@
Gaussian Splatting (GS) [kerbl_3d_2023, zhang2024gaussianimage], a recent method originally designed for 3D scene representation and rendering using volumetric Gaussian kernels—and conceptually related to earlier surface splatting techniques [Zwicker2001surfacesplatting], shares several similarities with SMoE. Both methods splat steered Gaussians into the pixel domain and use gradient descent to optimize Gaussian kernels for effective data representation. GS leverages localized block-rasterization to accelerate training, primarily targeting 3D applications. The key distinction between GS and SMoE lies in the GS-RBF regression model versus the SMoE normalized gating network for regression. SMoE regression can be seen as an extended, more powerful strategy compared to GS-RBF with normalization using the same kernel representation.
@@
First attempts to adopt the rasterized GS strategy to 2D image regression appeared in [zhang2024gaussianimage]. Building decisively on this foundation, our work is the first to integrate GS-inspired rasterization directly into the SMoE framework, combining GS’s computational speed with SMoE’s superior sparsity and edge-aware reconstruction. The purpose of this paper is to take advantage of both GS (fast) and SMoE (sparse and high-quality reconstruction) approaches to arrive at a fast and sparse high-quality SMoE regression method. To this end, we seek to adopt the rasterization approach of GS to optimize SMoE parameters, overcoming the critical bottleneck of slow, iterative gating network optimization and enabling drastic run-time improvements.
@@
  • We introduce R-SMoE, a rasterized SMoE training framework that leverages tile-based rasterization inspired by Gaussian Splatting. Compared to non-rasterized SMoE regression, R-SMoE accelerates model training and rendering times by orders of magnitude while maintaining high fidelity. Compared to “GaussianImage” [zhang2024gaussianimage], which employs GS-RBF rasterization, our approach requires significantly fewer computational resources for both training and rendering.
@@
Figure 1: Illustration of the edge-aware Steered Mixture of Experts (SMoE) model applied to image compression and denoising at 0.43 bits per pixel (bpp). Figure adapted from [fleig_edge-aware_2022].
@@ II Steered Mixture of Experts (SMoE)
The SMoE image model describes an edge-aware, parametric, continuous nonlinear regression function. Fig. 1 (from [fleig_edge-aware_2022]) illustrates the soft-gating concept of the kernel model for image compression and denoising tasks.
@@ II-A Theory
The Steered Mixture of Experts (SMoE) model extends traditional kernel regression methods, such as Radial Basis Function (RBF) [ghosh_overview_2001] networks used in Gaussian Splatting, by introducing a more sophisticated mechanism for image reconstruction. Although Gaussian Splatting also uses a “weighted sum of Gaussians,” which is conceptually similar to RBF regression, the interpretation differs: in RBF, the weights represent the importance of each Gaussian basis, whereas in Gaussian Splatting, the weights correspond to the color (appearance) of the Gaussian. Thus, while the regression form is similar, the modeling objectives are fundamentally different. RBF networks define the regression function $y_{p}(x)$ based on a weighted sum of $L$ (steered 2D Gaussian) kernels $K_{j}(x)$:
@@
The $\pi_{i}$ values provides additional weights to each kernel. Expert functions $m_{j}(x)$ can take various functional forms, including constant, linear, quadratic, and basis functions such as DCT and wavelet bases [blu2002wavelets]. For our work, we utilize simple constant experts $m_{j}(x)=m_{j}$, which have demonstrated excellent performance in previous studies [bochinski_regularized_2018, tok_mse_2018]. The irregularly shaped 2D softmax gating functions $w_{j}(x)$ are derived based on the position and parameters of the $L$ kernels, allowing for precise modeling of both smooth transitions and sharp edges in images.
@@ III-A Deep-Learning-Based Optimization
Recently, several methods have been proposed to replace traditional gradient descent (GD) optimization with deep-learning approaches. These methods train neural networks to directly predict kernel parameters. For instance, Fleig et al. [fleig_edge-aware_2022, fleig_steered_2023] trained an auto-encoder network to predict kernel parameters with a SMoE regressor according to Eq. (3) as the decoder. This results in drastically faster optimization, replacing the cumbersome GD process. However, these approaches are currently applicable for a maximum of 16 kernels on small blocks of 8x8 or 16x16 pixels. Currently, most approaches use a fixed number of kernels for each block, which provides limited adaptation to content. Additionally, because kernel parameter prediction is based on small, isolated blocks, independent predictions across different blocks can introduce blocking effects, where discontinuities or artifacts may appear between blocks.
@@ III-B Global Optimization
Global optimization [bochinski_regularized_2018, jongebloed_sparse_2022] represents the forefront of optimization strategies for kernel regression and SMoE models. This method optimizes the model by considering the entire set of kernels simultaneously, rather than treating them individually or in isolated groups. The core idea is to ensure that all kernels in the model contribute to the reconstruction of each pixel, thereby capturing the holistic structure of the image. This strategy is also common in RBF type regression [dale2007globalRBF]. Global optimization thus involves using all kernels to jointly reconstruct each pixel. Strategies exist to constrain the impact of kernels to the immediate neighborhood [verhack2018progressive] to reduce the complexity of reconstruction.
@@ III-C Rasterized optimization
Recently, rasterization [rasterization_laine_2011, rasterization_schotz_2022] has become an increasingly valuable tool for enhancing gradient descent (GD) optimization in image processing tasks. One notable example is the GS for 2D images (GaussianImage [zhang2024gaussianimage]), which leverages rasterization to boost the efficiency of GD optimization in Radial Basis Function (RBF) [ghosh_overview_2001] networks. Adopted from 3D GS [kerbl_3d_2023], this approach is similar to deep-learning-based optimization, being block-based in nature. Each image is divided into adjacent, non-overlapping blocks of 16x16 pixels. A very large number of 2D kernels are initialized over the image domain. A small subset of the kernels is identified as relevant for regressing a block prior to GD optimization, which allows fast and accurate GD optimization and block reconstruction. In this paper, we adapt the GS rasterized optimization strategy to SMoE regression, resulting in fast, sparse, and high-quality reconstruction.
@@ III-D Hybrid EM-Based Parallel Rendering
In addition to gradient-descent-based rasterized optimization, prior work by Avramelos et al. [Avramelos2020highlyparallelsmoe] introduced a block-parallel rendering method for SMoE models, designed for real-time applications. Their method relies on a two-phase pipeline: a global Expectation-Maximization (EM) optimization followed by block-level parallel rendering. During optimization, Gaussian kernel parameters—position, orientation, and scale—are learned globally using EM. The rendering stage then employs a fixed-radius search around each pixel to determine which kernels to evaluate, enabling efficient per-pixel computation.
@@ IV Rasterized SMoE
The pipeline for the proposed rasterized SMoE is illustrated in Fig. 3 (a). This pipeline mirrors that of [zhang2024gaussianimage], except that the RBF regression is replaced with the SMoE gating framework (c.f. Equ. (2) and (4)).
@@ IV-B Geometry shader
For each 16x16 pixel block, each Gaussian kernel of the set $\mathcal{K}$ is given a confidence ellipse corresponding to a 99% confidence interval of its 2D Gaussian distribution. This ellipse approximates the region in which the kernel has significant influence, with its major and minor axes derived from the kernel’s covariance matrix. For the mathematical intricacies—including axis derivation via eigen-decomposition—see [genz2009computation].
@@ V Rasterized SMoE for Denoising
While previous SMoE methods have employed block-based representations for denoising, either through gradient descent (GD) [ozkan_steered_2023] or deep learning (DL) [fleig_steered_2023] optimization, these approaches are limited by their block constraints. R-SMoE promises a more efficient and reliable representation, combining the advantages of global representation with the flexibility of local processing. R-SMoE does not confine the influence of kernels to a small block, allowing each kernel to contribute beyond its immediate neighborhood.
One significant challenge when applying GS for 2D images and SMoE models to denoising tasks is that, when applied to smaller blocks of 16x16 pixels, these models attempt to accurately reconstruct not only the signal but also part of the noise. A promising strategy is then to employ overlapping GS blocks or R-SMoE blocks and reconstruct the pixels using multiple models in a multi-model approach [ozkan_steered_2023]. This yields a natural extension: multi-model inference in R-SMoE (MM-RSMoE), which we explore for denoising.